OpenStudy (anonymous):

Karen is 13 years older than Juan. The sum of their ages is 55. How old is Juan? juan is 13

OpenStudy (anonymous):

x+x+13 =55

OpenStudy (anonymous):

Let Juan's age = x Karen's age = x + 13 If the SUM of their ages is 55, then x + (x+13) = 55. Now, you can finish the problem.

OpenStudy (anonymous):

42

OpenStudy (anonymous):

No, Juan's age was represented by x. x cannot be 42. If Juan is 42, then Karen's age will be 55, and their sum is 97. Solve the equation I wrote for you.

OpenStudy (anonymous):

Can you solve that equation?

OpenStudy (anonymous):

I am doing it now, one sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

juan is 21

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

Yes. Congratulations.

OpenStudy (anonymous):

Yay lol.... thank you!

OpenStudy (anonymous):

welcome.